OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If line is undefined, state so.... (5,3) and (8,-4)

OpenStudy (lalaly):

slope = y2-y1 / x2-x1 =-4-3 / 8-5= -7/3

OpenStudy (anonymous):

Thank you so much :)